An object of mass 4kg moves round a circle of radius 6m with a constant speed of 12m/s. Calculate

(1) the angular velocity and the force towards the center


1
Expert's answer
2021-02-11T17:11:07-0500

"\omega=\frac vr,~" "\omega=\frac{12}{6}=2~\frac{\text{rad}}{\text{s}}."


"F=m\omega^2 r,~" "F=4\cdot2^2\cdot 6=96~\text{N}."